OpenStudy (anonymous):

Can someone help me w. the steps to solve (x)log(2)=(x+1)log(5) ??

OpenStudy (anonymous):

treat everything but \(x\) as a constant distribute first \[x\log(2)=x\log(5)+\log(5)\]

OpenStudy (anonymous):

Apparently I'm suppose to convert it into an exponential equation but idk where to start.

OpenStudy (anonymous):

put all \(x\)'s on one side \[x\log(2)-x\log(5)=\log(5)\] factor out an \(x\) from the left \[x(\log(2)-\log(5)=\log(5)0\] then divide \[x=\frac{\log(5)}{\log(2)-\log(5)}\]

OpenStudy (anonymous):

ty